# The PLATO Mission

### v1

@@ 4.1.2 Ultra-precise seismic stellar characterisation with Kepler
The detection of solar-like oscillations, the quality of the measured seismic parameters, and the accuracy with which seismic stellar characterisation is obtained all depend on the SNR, the cadence, and the observation baseline for each star. The SNR depends not only on the instrumental observational noise and the apparent magnitude of the star but also on the type of star. Hence, depending on the quality of the observations and the type of star, we have different levels of stellar characterisation. PLATO is designed and optimally suited to perform asteroseismology of sun-like stars. However, we learned from the Kepler data that, for K dwarfs and M-type stars, either the oscillation amplitudes are too low for detection or the stars do not oscillate at all [263, 262]. Even for F to G-type dwarfs, it is thought that the impact of strong surface magnetic activity on envelope turbulent convection can cause a significant decrease of the oscillation amplitudes, which can hamper their detection [120, 69, 174, 271]. Some other causes of non-detection are also possible, such as metallicity effects and binarity [267, 170, 205]. It is not yet clear whether TESS in its 20 second cadence mode with its limited photometric quality can lead to advances on these topics, but PLATO with its 25 second cadence mode and high photometric quality will bring new insights. On the hot side, late F-type stars have larger mode linewidths than G-type stars, causing larger uncertainties in their seismic measurements. These translate into larger statistical uncertainties for their stellar parameters [14, 86]. However, those stars with their large radii (larger than 1.2 solar radius) and high temperatures are not the main priority for PLATO’s exoplanet search and characterisation. Actually the P1 sample is truncated at temperatures on the hot side, such that one still encounters solar-like oscillators. We therefore focus hereafter on the case of late F, G, and early K-type dwarfs, and subgiants.
@@
When the SNR is low and/or the observation time duration is short (one month or so), the solar-like oscillations are detected mainly as a power excess in a power spectrum. One then takes the regularity of the frequency pattern to determine a first average seismic quantity, namely the mean large frequency separation, , that scales with the square root of the mean stellar density. Several techniques also allow the determination of the frequency at maximum oscillation power, , in a power spectrum. This frequency is related to the surface gravity of the star and to some lesser extent to the effective temperature. Given the effective temperature, these two quantities give direct access to the mass and radius of the star [143, see also the quoted reviews above and references therein]. Hence, uncertainties of the resulting mass and radius determination directly depend on the uncertainties of , , and . For a sample of about 500 Kepler stars observed in short cadence mode over a period of nearly a month, the quoted average median statistical uncertainties are 1.5 – 2% for and 4% – 5.8% for . For rather high SNR [310] these measured seismic observables and their uncertainty lead to a median relative uncertainty of the order of 5.5% and 10% in radius and mass, respectively [69]. Further Kepler data have shown that one must take into account the correlation between and due to stellar evolution. This correlation was empirically derived from Kepler data [287, 155]. Taking it into account in the modelling decreases significantly the statistical uncertainties in masses and radii [278].
@@
A more precise characterisation, including age determination, is obtained when one infers the properties of a star by means of a fitting technique. Usually a grid-based approach is adopted, where one selects the best stellar models from a large grid in such a way that the oscillations predicted for these models satisfy the observational constraints to within a defined matching criterion. Those best-fitting models are often referred to as “optimal (stellar) models” and their properties are attributed to the studied star. The interval of values of each property, such as mass, radius or age, spanned by the set of optimal models is then used to provide an estimate of the uncertainty for that property. Those approaches require that the observational constraints include not only the average seismic parameters but also classical observables such as the effective temperature and the metallicity of the star. The luminosity can be used as an additional constraint or to cross-check as validation after the modelling.
Several pipelines have been developed for asteroseismic modelling, each with the same goal but differing in various aspects. The inter-comparison of their results shows good agreement if all other aspects, such as input data and stellar models, are kept the same. One of the most popular techniques searches the optimal stellar models in a pre-computed grid and is therefore referred to as grid-based modelling (GBM) [e.g. 282, 261, 258, 187, 112, 7, and several others]. Because it is easy to use and relatively fast, the GBM approach is favoured for the PLATO baseline stellar modelling pipeline.
The use of seismic average parameters together with a GBM approach leads to a significant reduction of the uncertainties in the overall characterisation of a set of stars compared to those provided by classical modelling methods only, even when including interferometric constraints, as the age uncertainties can still reach more than 50 % in that case [204, 40]. With statistical uncertainties of 2% for and 4% for , the first GBM determinations for 500 stars with seismic one-month observations yielded median relative uncertainties of the order of 2.8%, 2.2%, 5.4%, and 25% in mean density, radius, mass, and age when spectroscopic effective temperatures are adopted [71]. A number of seismic modelling efforts characterised sets of Kepler stars, including planetary hosts. This led to homogeneously derived stellar properties based on the average seismic parameters [157, 199, 60, 204, 206, 24].
A homogeneous catalogue consisting of 415 stars taken from the Chaplin et al. sample is available, where the seismic analyses were based on long observation baselines up to 1050 days and spectroscopic and metallicity were used ([278]). This provided average median precisions of 1.7% and 4%, respectively, for and . A GBM stellar characterisation yielded median statistical errors of about 2.3%, 4.2%, and 15% for radius, mass, and age, respectively. We note that the faintest stars in PLATO’s P1 sample as well as a number of the brightest stars in the P5 sample are likely to be in this situation, meaning that for their mass, we are already within the PLATO requirements as far as precision is concerned. The PLATO age-dating requirement, however, still needs to be scrutinised with better models and modelling methodology, which is under current development.
@@
While acceptable results are obtained by using the average seismic quantities, the most precise stellar characterisation is obtained when a detailed measurement of the frequencies of individual oscillation modes is possible. This occurs for the brightest stars observed over a long temporal baseline. In such cases, individual mode frequencies can be fitted with methods such as the popular “peak-bagging” technique [e.g. 13, 139, 88]. In this way the Kepler mission has provided precise seismic and stellar parameters for about 150 dwarfs and subgiant stars, including planetary hosts, using individual frequencies of oscillation modes and various frequency-difference combinations. Combined with optimisation procedures such as a GBM approach, this allowed for a very precise determination of the properties of these stars. The precision obtained with detailed modelling are somewhat dependent on the chosen quantity to be reproduced (frequencies or frequency combinations), as well as the optimisation algorithm and the density of the stellar evolution model grids in terms of the free stellar parameters involved in the fitting procedure.
At present, two major catalogues of stars exist from the viewpoint of containing the best homogeneously derived oscillation frequencies, stellar masses, radii, and ages: the so-called Kages sample and the Legacy sample. The Kages sample includes 35 exoplanet host stars, which were observed over the whole nominal Kepler mission, resulting in oscillation frequency uncertainties at the level of 0.3 Hz (best cases 0.1 Hz) in the vicinity of for dipolar modes [96]. [282] used combinations of the individual frequencies and a GBM approach to derive stellar properties with median statistical uncertainties of 1.7% (density), 1.2% (radius), 3.3% (mass), and 14% (age). The Legacy sample includes 66 dwarfs of low-mass with short-cadence Kepler observations spanning at least 1 year and up to 4 years. With a white noise level ranging between 0.1 and 8 Hz (5.3 – 47 ), frequency uncertainties at for dipolar modes are found in the range Hz depending on the star [198]. The first stellar modelling based on these high quality data was carried out by seven different teams. All teams used their own pipelines, which differ by their optimisation methods, evolution and oscillation codes, input physics assumptions, and seismic diagnostics (individual frequencies or frequency combinations). Quoted average uncertainties range from 0.5% to 2.6% in density, 1.3% to 4.2% in radius, 2.3% to 4.5% in mass, and 6.7% to 20% in age [283].
Following these pioneering ensemble modelling studies, a number of theoretical studies flourished using the Legacy sample or a subset of this sample that have confirmed the ability of asteroseismology to provide stellar masses and ages with precisions of a few tens of a percent [e.g. 90, 35, 237, 111, 112, 7]. As an example, [234] found uncertainties of typically 10-20% for the ages and masses at the level of 2-4%, while lower than 0.01 dex for surface gravities for the Kepler Legacy sample.
@@
For cool stars, the TESS survey is more adapted to M-dwarf observations. Yet, in its short cadence mode, TESS provides data enabling the detection of solar-like oscillations and characterisation of very bright () main-sequence stars [73, 159, 141]. TESS is better suited for seismic studies of subgiants due to their larger oscillation amplitudes compared to those of main-sequence stars [72, 212]. However, the short duration of the TESS observations results in the detection of only a few modes around . Due to their evolved structure the subgiant frequency spectra include a few mixed modes. The data analysis process and characterisation procedure based on mixed modes has to be more sophisticated but can lead to an extremely precise characterisation in mass and age [see for instance 23, 232, 190, 235].
Summarizing at this stage, the Kages and Legacy stars together with a few other individual seismically studied stars from CoRoT, Kepler/K2 and TESS represent today’s set of most precisely characterised field stars. The addition of asteroseismology to stellar modelling permits to reach levels of 10% precision in stellar ages at least for stars similar to the Sun [see for the case of Kepler 93 43]. This complies with the requirements for the PLATO mission as far as precision is concerned. This was further confirmed by a PLATO dedicated hare-and-hound exercises [94, see below]. Together with the Sun, those stars are now being used to calibrate gyrochronology relations, which can then be used to age-dating other stars without detected oscillation modes. Those stellar samples also serve as reference stars to validate other techniques capable of constraining stellar properties, such as the derivation of from photometric variability due to surface granulation ([54] and references therein).
@@ 4.2 Precision versus accuracy
Several studies using simulations have shown that, given the same input set of information, the various data-analysis and optimisation methods have reached such a high level of sophistication that they do not introduce large biases and yield similar results. In contrast, even restricting the case to low-mass main-sequence and subgiant stars as we do here, important sources of systematic errors result from our poor knowledge of various physical processes that affect the stellar structure and evolution, as well as the oscillation frequencies. This is particularly true for the age, which of all stellar properties, is by far the most challenging to determine accurately. In most cases, stellar ages can only be determined through stellar modelling and their accuracy therefore strongly depends on the degree of reliability of the available stellar models [e.g., 76]. Particularly critical (and uncertain) are physical processes that mix the chemical elements in the stellar cores, affecting the relation between the age and the composition profile and, thus, the seismically inferred age.
@@
Accuracy tests: independent additional observational information. For Kepler stars using interferometric or astrometric observations as external constraints, scaling relations have been reported to be accurate at the level of in density, 2 – 5 in radius, and about in masses for dwarfs and subgiants [e.g. 83, 158, 321]. Taking into account several different sources of systematic error in a GBM approach, [278] quoted a systematic error of the order of 1.2 in radius, 3 in mass and 12 in age with a total combined error of approximately 2.6 in radius, 5.1 in mass, and 19 in age [see Tab.2 in 278]. Those studies also emphasised the importance of having accurate and precise classical parameters to exploit the full potential of seismology. Getting such classical parameters requires the development of specific and sophisticated pipelines based on as accurate model atmospheres as possible [223, 125]. This by itself represents a whole branch of stellar astrophysics, which contributes to properly characterise stars.
@@
Accuracy tests: “differential” studies. Another commonly used way of at least partially assessing the accuracy is rather differential than absolute: it consists of considering the dispersion of the results provided by different pipelines (which differ in many aspects of input physics) or the results coming from a single pipeline but varying the input physics of the stellar models. In this context, several studies have revisited the characterisation of the Kages and Kepler Legacy stars. It was found that significant systematic differences for the mean values of the mass and the age occur. The dispersion in medium age can range between 15% and 33%. This is attributed to different options about the optimisation strategy but mostly to different options of the stellar modelling [34, 35, 90, 111, 237, 236, 238]. The most studied Legacy stars are the brightest (G1.5V + G3V) solar analogue components A and B of the multiple system 16 Cyg (with apparent Kepler magnitudes and , respectively). Their masses are precisely determined at the level of 3% or below. Most studies passed the accuracy test of both stars having the same age – as expected for a binary system – within the quoted uncertainties (reported to be in the range 2 – 6% depending on the study) [e.g. 283, 90, 34, 35, 55, 309, 31]. However, examining the mean values derived by those various studies tells us that the seismically derived age of 16 Cyg A covers a range of Gyr depending on the assumed input physics. [111] modelled both stars varying one input physical ingredient at a time and found ages in the range 6.4 () to 7.5 () Gyr. This represents an age dispersion up to about 16% when one takes 6.95 Gyr as a reference mean value for 16 Cyg A. All these above dispersion estimates are larger than the uncertainties provided by each individual pipeline and currently remain larger than the 10 % accuracy for such a type of star as required for the PLATO mission.
@@
Although the various sources of systematic errors due to stellar modelling cannot be detailed here, it is clear that improvements in stellar modelling are still necessary in the coming years in order to reach the PLATO age-dating requirement. This is within reach from intense on-going theoretical work within the PMC addressing the main inaccuracy issues. Real advances in this activity will come from the confrontation of the updated modelling with observations of a sample of well-characterised stars (i.e., with the highest possible precision and accuracy, mainly via seismology).
@@ 4.3 Detecting solar-like oscillations with PLATO
The requirements of the PLATO mission are extremely challenging, especially for the stellar age (e.g. 10% accuracy for a reference star with mag with – that is a solar analogue slightly hotter than our Sun). In the following, the assessment of the seismic performance of PLATO is discussed at two levels: the detection of oscillations as an excess of power in a power spectrum that comes with the measurement of the two averaged seismic parameters and the detection of oscillation modes for which individual frequencies will be measured. We will consider the performance for the sample P1–P2 on the one hand and the sample P5 on the other hand. The estimates are based on what Kepler data taught us (for details see, Goupil et al. in prep. ).
The P1–P2 sample is deliberately designed as composed of stars with a noise level (random and systematic residual, non-stellar) below at magnitude . It is therefore made up of a large number of targets for which one expects the detection of solar-like oscillations in the majority of cases (K-stars remain an open issue) as well as the measurement of individual mode frequencies with a precision high enough to provide high quality seismic masses, radii, and ages. We expect a considerable improvement in the quality of stellar modelling in terms of accuracy thanks to the P1–P2 seismic sample.

### v2

@@ 4.1.2 Ultra-precise seismic stellar characterisation
The detection of solar-like oscillations, the quality of the measured seismic parameters, and the accuracy with which seismic stellar characterisation is obtained all depend on the SNR, the cadence, and the observation baseline for each star. The SNR depends not only on the instrumental observational noise and the apparent magnitude of the star but also on the type of star. Hence, depending on the quality of the observations and the type of star, we have different levels of stellar characterisation. PLATO is designed and optimally suited to perform asteroseismology of sun-like stars. However, we learned from the Kepler data that, for K dwarfs and M-type stars, either the oscillation amplitudes are too low for detection or the stars do not oscillate at all [286, 285]. Even for F to G-type dwarfs, it is thought that the impact of strong surface magnetic activity on envelope turbulent convection can cause a significant decrease of the oscillation amplitudes, which can hamper their detection [131, 75, 186, 294]. Some other causes of non-detection are also possible, such as metallicity effects and binarity [290, 182, 226]. On the hot side, late F-type stars have larger mode linewidths than G-type stars, causing larger uncertainties in their seismic measurements. These translate into larger statistical uncertainties for their stellar parameters [15, 91]. However, those stars with their large radii (larger than 1.2 solar radius) and high temperatures are not the main priority for PLATO’s exoplanet search and characterisation. Actually the P1 sample is truncated at temperatures on the hot side, such that one still encounters solar-like oscillators. We therefore focus hereafter on the case of late F, G, and early K-type dwarfs, and subgiants.
@@
When the SNR is low and/or the observation time duration is short (one month or so), the solar-like oscillations are detected mainly as a power excess in a power spectrum. One then takes the regularity of the frequency pattern to determine a first average seismic quantity, namely the mean large frequency separation, , that scales with the square root of the mean stellar density. Several techniques also allow the determination of the frequency at maximum oscillation power, , in a power spectrum. This frequency is related to the surface gravity of the star and to some lesser extent to the effective temperature. Given the effective temperature, these two quantities give direct access to the mass and radius of the star [see 156, and references therein]. Hence, uncertainties of the resulting mass and radius determination directly depend on the uncertainties of , , and . For a sample of about 500 Kepler stars observed in short cadence mode over a period of nearly a month, the quoted average median statistical uncertainties are 1.5 – 2% for and 4% – 5.8% for . For rather high SNR [334] these measured seismic observables and their uncertainty lead to a median relative uncertainty of the order of 5.5% and 10% in radius and mass, respectively [75]. As shown by [142], most of the stars within the PLATO P1 sample and a large number of stars in the P5 sample will benefit from such uncertainties.
@@
A more precise characterisation, including age determination, is obtained when one infers the properties of a star by means of a fitting technique. Usually a grid-based approach is adopted, where one selects the best stellar models from a large grid in such a way that the oscillations predicted for these models satisfy the observational constraints to within a defined matching criterion. In that spirit, several pipelines have been developed for asteroseismic modelling, each with the same goal but differing in various aspects. The inter-comparison of their results shows good agreement if all other aspects, such as input data and stellar models, are kept the same. One of the most popular techniques searches the optimal stellar models in a pre-computed grid and is therefore referred to as grid-based modelling (GBM) [e.g. 305, 284, 281, 204, 121, 8, and several others].
Because it is easy to use and relatively fast, the GBM approach is favoured for the PLATO baseline stellar modelling pipeline. This is motivated by the Kepler experience, which shows that for long duration observations (up to 1050 days), we obtain average median precisions of 1.7% and 4%, respectively, for and and subsequently median statistical errors of about 2.3% for radius, 4.2% for mass, and 15% for age. This result has been obtained using a GBM stellar characterisation [302] and implies that the faintest stars in PLATO’s P1 sample as well as a number of the brightest stars in the P5 sample are likely to be in this situation, meaning that for their mass, we are already within the PLATO requirements as far as precision is concerned. The PLATO age-dating requirement, however, still needs to be scrutinised with better models and modelling methodology, which is under current development.
@@
While acceptable results are obtained by using the average seismic quantities, the most precise stellar characterisation is obtained when a detailed measurement of the frequencies of individual oscillation modes is possible. This occurs for the brightest stars observed over a long temporal baseline. In such cases, individual mode frequencies can be fitted with methods such as the popular “peak-bagging” technique [e.g. 14, 152, 93]. Combined with optimisation procedures such as a GBM approach, this allowed for a very precise determination of the properties of these stars. The precision obtained with detailed modelling are somewhat dependent on the chosen quantity to be reproduced (frequencies or frequency combinations), as well as the optimisation algorithm and the density of the stellar evolution model grids in terms of the free stellar parameters involved in the fitting procedure.
At present, two major catalogues of stars exist from the viewpoint of containing the best homogeneously derived oscillation frequencies, stellar masses, radii, and ages: the so-called Kages sample and the Legacy sample. The Kages sample includes 35 exoplanet host stars, which were observed over the whole nominal Kepler mission, resulting in oscillation frequency uncertainties at the level of 0.3 Hz (best cases 0.1 Hz) in the vicinity of for dipolar modes [101]. [305] used combinations of the individual frequencies and a GBM approach to derive stellar properties with median statistical uncertainties of 1.7% (density), 1.2% (radius), 3.3% (mass), and 14% (age). The Legacy sample includes 66 dwarfs of low-mass with short-cadence Kepler observations spanning at least 1 year and up to 4 years. With a white noise level ranging between 0.1 and 8 Hz (5.3 – 47 ), frequency uncertainties at for dipolar modes are found in the range Hz depending on the star [218]. Quoted average uncertainties range from 0.5% to 2.6% in density, 1.3% to 4.2% in radius, 2.3% to 4.5% in mass, and 6.7% to 20% in age [306].
@@
Summarizing at this stage, the Kages and Legacy stars together with a few other individual seismically studied stars from CoRoT, Kepler/K2 and TESS represent today’s set of most precisely characterised field stars. The addition of asteroseismology to stellar modelling permits to reach levels of 10% precision in stellar ages at least for stars similar to the Sun [see for the case of Kepler 93 42]. This complies with the requirements for the PLATO mission as far as precision is concerned. This was further confirmed by a PLATO dedicated hare-and-hound exercises [99, see below]. Together with the Sun, those stars are now being used to calibrate gyrochronology relations, which can then be used to age-dating other stars without detected oscillation modes. Those stellar samples also serve as reference stars to validate other techniques capable of constraining stellar properties, such as the derivation of from photometric variability due to surface granulation ([59] and references therein).
@@ 4.2 Precision versus accuracy
A number of studies have used several fitting methods to derive the stellar mass, radius and age of oscillating solar-type stars with an internal structure similar to that of the Sun (e.g. [203], [305], [279], [99]). They found that, given the same set of input information (same grid of stellar models, same number of free parameters to adjust, same choice of seismic diagnostics, …), the different data analysis and optimisation methods give similar results (with uncertainties well below PLATO requirements) and introduce much smaller biases than those induced by systematic errors due to the choice of seismic diagnostics, to the adopted number of free parameters or, more importantly, to the physical description of the stars.
In contrast, even restricting the case to low-mass main-sequence and subgiant stars as we do here, important sources of systematic errors result from our poor knowledge of various physical processes that affect the stellar structure and evolution, as well as the oscillation frequencies. This is particularly true for the age, which of all stellar properties, is by far the most challenging to determine accurately. In most cases, stellar ages can only be determined through stellar modelling and their accuracy therefore strongly depends on the degree of reliability of the available stellar models [e.g., 81]. Particularly critical (and uncertain) are physical processes that mix the chemical elements in the stellar cores, affecting the relation between the age and the composition profile and, thus, the seismically inferred age.
@@
Accuracy tests: independent additional observational information. For Kepler stars using interferometric or astrometric observations as external constraints, scaling relations have been reported to be accurate at the level of in density, 2 – 5 in radius, and about in masses for dwarfs and subgiants [e.g. 88, 170, 345]. Taking into account several different sources of systematic error in a GBM approach, [302] quoted a systematic error of the order of 1.2 in radius, 3 in mass and 12 in age with a total combined error of approximately 2.6 in radius, 5.1 in mass, and 19 in age [see Tab.2 in 302]. Those studies also emphasised the importance of having accurate and precise classical parameters to exploit the full potential of seismology. Getting such classical parameters requires the development of specific and sophisticated pipelines based on as accurate model atmospheres as possible [246]. This by itself represents a whole branch of stellar astrophysics, which contributes to properly characterise stars. We note that the determination of the classical parameters is an integral part of the mission, as developing a specific pipeline that makes use of various types of observations to tightly constrain these quantities is one of the ground-segment activities ([135], Orlander et al., 2024 submitted). In parallel, there are efforts to set up collaborations with large-scale surveys (e.g. 4MOST; Guiglion et al [148]) in order to optimise the homogeneity and completeness of the preparatory spectra available for the PLATO targets. It will allow to meaningfully examine the correlations of planet properties and their occurrence rates with, for instance, the metallicity of the stellar host.
@@
Accuracy tests: “differential” studies. Another commonly used way of at least partially assessing the accuracy is rather differential than absolute: it consists of considering the dispersion of the results provided by different pipelines (which differ in many aspects of input physics) or the results coming from a single pipeline but varying the input physics of the stellar models. In this context, several studies have revisited the characterisation of the Kages and Kepler Legacy stars. It was found that significant systematic differences for the mean values of the mass and the age occur. The dispersion in medium age can range between 15% and 33%. This is attributed to different options about the optimisation strategy but mostly to different options of the stellar modelling [34, 35, 95, 120, 257, 256, 258]. The most studied Legacy stars are the brightest (G1.5V + G3V) solar analogue components A and B of the multiple system 16 Cyg (with apparent Kepler magnitudes and , respectively). Their masses are precisely determined at the level of 3% or below. Most studies passed the accuracy test of both stars having the same age – as expected for a binary system – within the quoted uncertainties (reported to be in the range 2 – 6% depending on the study) [e.g. 306, 95, 34, 35, 60, 333, 31]. However, examining the mean values derived by those various studies tells us that the seismically derived age of 16 Cyg A covers a range of Gyr depending on the assumed input physics. [120] modelled both stars varying one input physical ingredient at a time and found ages in the range 6.4 () to 7.5 () Gyr. This represents an age dispersion up to about 16% when one takes 6.95 Gyr as a reference mean value for 16 Cyg A. All these above dispersion estimates are larger than the uncertainties provided by each individual pipeline and currently remain larger than the 10 % accuracy for such a type of star as required for the PLATO mission.
@@
Although the various sources of systematic errors due to stellar modelling cannot be detailed here, it is clear that improvements in stellar modelling are still necessary in the coming years in order to reach the PLATO age-dating requirement. This is within reach from intense on-going theoretical work within the PMC addressing the main inaccuracy issues. Real advances in this activity will come from the confrontation of the updated modelling by the time of the PLATO commissioning with observations during the first long pointing of a sample of well-characterised stars (i.e., with the highest possible precision and accuracy, mainly via seismology).
We point out that the above discussion on systematic uncertainties only concern solar-like stars in the core science case of PLATO. The seismology of stars born with a convective core involve other uncertainties, particularly for stars that are not subject to magnetic braking due to lack of a convective envelope. We refer the reader to the reviews by Aerts [5] and Aerts and Tkachenko [6] for modelling procedures of such objects and the accompanying uncertainties. These stars are part of PLATO’s Complementary Science program.
@@ 4.3 Detecting solar-like oscillations with PLATO
The requirements of the PLATO mission are extremely challenging, especially for the stellar age (e.g. 10% accuracy for a reference star with mag with – that is a solar analogue slightly hotter than our Sun). In the following, the assessment of the seismic performance of PLATO is discussed at two levels: the detection of oscillations as an excess of power in a power spectrum that comes with the measurement of the two averaged seismic parameters and the detection of oscillation modes for which individual frequencies will be measured. We will consider the performance for the sample P1–P2 on the one hand and the sample P5 on the other hand. The estimates given below are based on what Kepler data taught us (for details see, [142] ).
The P1–P2 sample is deliberately designed to be composed of stars with a noise level (random and systematic residual, non-stellar) below at magnitude . It is therefore made up of a large number of targets for which one expects the detection of solar-like oscillations in the majority of cases (K-stars remain an open issue) as well as the measurement of individual mode frequencies with a precision high enough to provide high quality seismic masses, radii, and ages. We expect a considerable improvement in the quality of stellar modelling in terms of accuracy thanks to the P1–P2 seismic sample.